OpenStudy (anonymous):

I just an explanation of the second part, Please!! a) At what level of production will the marginal cost be minimized? b) At what level of production will the average cost A(x)=C(x)/x be minimized? C(x)= x^(3)-12x^(2)+48x+98

OpenStudy (anonymous):

just plug in your number for x

OpenStudy (bahrom7893):

Well you just take A(x), which is: C(x)/x = x^2 - 12x + 48 + 98x^(-1) Take the derivative of A(x): A'(x) = 2x - 12 - 98x^(-2)

OpenStudy (bahrom7893):

Set that equal to 0, solve for x.

OpenStudy (anonymous):

I still don't get the answer.

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

ok.. let's see

OpenStudy (anonymous):

By the principle, if Average Cost is Minimum, <=> Average Cost = Marginal Cost

OpenStudy (anonymous):

You've already had Marginal Cost in question a)

OpenStudy (anonymous):

I think so, lol! I have 3(x-4)^2 that's for the a)

OpenStudy (anonymous):

How do get Marginal Cost?

OpenStudy (anonymous):

I follow the examples, sorry!

OpenStudy (anonymous):

I believe that is kind of factoring..

OpenStudy (anonymous):

for the first two numbers.

OpenStudy (anonymous):

No, Marginal Cost is C' (x)

OpenStudy (anonymous):

Ohh!!

OpenStudy (anonymous):

ok, I got 3x^2-24x+48 right!

OpenStudy (anonymous):

Okie, it's just another way of minimum value :)

OpenStudy (anonymous):

Simplified: -> Marginal Cost C'(x) = x² - 8x + 16

OpenStudy (anonymous):

Now set Marginal Cost = Average Cost

OpenStudy (anonymous):

Ok, that's is my problem..

OpenStudy (anonymous):

How you got the the marginal Cost C'(x)? I have it different way.

OpenStudy (anonymous):

That's the same thing as this 3(x-4)^2?

OpenStudy (anonymous):

Your way's correct!

OpenStudy (anonymous):

Ok, Uff!!!

OpenStudy (anonymous):

Ok, hold on then!! lol!

OpenStudy (anonymous):

How do i set it up? Like this

OpenStudy (anonymous):

Like "this" ???

OpenStudy (anonymous):

2x^3-12x^2-98=0

OpenStudy (anonymous):

Where does it come from?

OpenStudy (anonymous):

I gues that the last part, f(x)

OpenStudy (anonymous):

You see what i mean, is the way that the explain i don't get it.

OpenStudy (anonymous):

It is an online class.

OpenStudy (anonymous):

It's extremely confused to solve from the ... middle of a problem !!!

OpenStudy (anonymous):

Now I think I must back track from the beginning :/

OpenStudy (anonymous):

Ok. How do i set up the Marginal Cost = Average Cost??

OpenStudy (anonymous):

So which one do you think is Marginal Cost?

OpenStudy (anonymous):

Ok, either yours or mine! 3(x-4)^2

OpenStudy (anonymous):

It's the same, but my equation is already expanded and simplify, so use it!

OpenStudy (anonymous):

Ok nice!

OpenStudy (anonymous):

Which one is Average Cost?

OpenStudy (anonymous):

Write them here with "= " in between :)

OpenStudy (anonymous):

x=16?

OpenStudy (anonymous):

Please show your work, I don't have the talent to read people's mind!

OpenStudy (anonymous):

How? Not sure how to do it! :(

OpenStudy (anonymous):

What do you mean? Where do you get x = 16???

OpenStudy (anonymous):

So please do NOT guess, as I don't guess and trying to help you out here!

OpenStudy (anonymous):

I know that's why i am asking you, because i don't know how to do it.. Sorry!

OpenStudy (anonymous):

Marginal Cost = Average Cost

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

3x^2-24x+48=3(x-4)^2

OpenStudy (anonymous):

This is just Marginal Cost equation! Plug the Average equation to the right side, instead!

OpenStudy (anonymous):

No way I gave up!

OpenStudy (anonymous):

Did you see the other solver already posted above?

OpenStudy (anonymous):

x^2-12x+48+98=0

OpenStudy (anonymous):

C(x)/x = x^2 - 12x + 48 + 98/x

OpenStudy (anonymous):

It's Average Cost function!

OpenStudy (anonymous):

Yes I have that part.

OpenStudy (anonymous):

i forgot the x sorry!

OpenStudy (anonymous):

Then set Marginal Cost = Average Cost

OpenStudy (anonymous):

3x^2-24x+48=2x^3-12x^2-98/x .

OpenStudy (anonymous):

C(x)/x = x^2 - 12x + 48 + 98/x

OpenStudy (anonymous):

@rocal2 What's so difficult to copy 2 equations down???

OpenStudy (anonymous):

It is not difficult, believe me this question is the only one that it is drivie me crazy..

OpenStudy (anonymous):

I am not sure what i am doing. That's why i am here!

OpenStudy (anonymous):

I keep showing you which one is Average function, and you keep copying down incorrect ???

OpenStudy (anonymous):

I am telling that i don't how to solve it...

OpenStudy (anonymous):

But that's ok, thanks for trying, and i am sorry to waste your time, I didn't meant to!!

OpenStudy (anonymous):

I will go to my school and find someone who can explaine me... Thank you so much!!